# Design and Evaluation of SEANet: a Software-defined Networking Platform for the Internet of Underwater Things

### v1

@@ Abstract
Investigating and safeguarding our oceans is vital for a host of applications and tasks, including combating climate change, ensuring the integrity of subsea infrastructures, and for coastal protection. Achieving these essential functions depends on the deployment of cost-effective, versatile underwater sensor networks that can efficiently collect and transmit data to land. However, the success of such networks is currently hindered by the significant limitations of existing underwater modems, which limits their operational use to a narrow range of applications. This paper presents and evaluates the performance of the SEANet software-defined networking platform, for the Internet of Underwater Things (IoUT), addressing the limitations of existing underwater communication technologies. It presents the development and comprehensive testing of an adaptable, high-data-rate, and integration-friendly underwater platform that reconfigures in real-time to meet the demands of various marine applications. With an acoustic front end, the platform significantly outperforms conventional modems, achieving more than double the data rate at 150 kbit/s. Experiments conducted in oceanic conditions demonstrate its capabilities in channel characterization, OFDM link establishment, and compatibility with the JANUS communication standard. Our platform advances the IoUT by providing a versatile, scalable solution that can incorporate multiple physical layers and support an array of tasks, making it pivotal for real-time ocean data analysis and the expansion of ocean-related digital applications.
@@
Acoustic modems have been shown to be viable solutions for long-range communications, as radio and optical technologies cannot be used by devices more than a few meters apart underwater [2]. Even if acoustic modems have been available for decades, current devices are beset by a host of limitations, including low data rates, cumbersome deployment procedures, and inflexible architectures. As such, many of the desired applications of the IoUT cannot be implemented with available technology, which motivates research and development of new devices that overcome current restrictions.
@@ III-A2 Converter Module
The converter module interfaces the main module with the rest of the modules. It has analog-to-digital (ADC) and digital-to-analog (DAC) data converters that provide analog interfaces to and from communication modules. In addition, it has voltage regulators for the main module, clock sources, and digital I/O ports. Since the platform is based on a direct conversion architecture and relies on digital mixers, to support wide range of frequencies used for underwater communications, fast data converters are chosen.
@@ III-A4 Power Module
The power module is an interface to the battery unit and powers each module with different voltages using onboard voltage regulators. The module includes filters to reduce power supply noise and protection circuitry that continuously monitors the voltage and current levels delivered to the platform.
@@ III-A5 Signal Module
The signal module routes digital and analog signals between the converter and communication modules and the transducer. As in the power module, each signal is accessible from any board in the stack. The signal module mounts reconfigurable board edge connectors to provide flexibility, e.g., it is possible to have more than one converter module attached to it. The signal module also has a connector for the transducer interface. This connection is driven by a duplexer circuit whose inputs can be switched between the power amplifier and preamplifier channels to allow a single transducer to be used for bidirectional communication. Even though transmit and receive chains can operate simultaneously, a half-duplex operation is implemented mainly due to the high cost of high-bandwidth acoustic transducers.
@@ IV Prototyping
This section describes the realization of the design in Section III. We start with the hardware implementation, followed by the software components. Finally, we discuss the realization of the mechanical structure.
@@ IV-A Hardware Implementation
In this part, we describe the key electronic components that we used to realize the desired design features.
@@ IV-A1 Main Module
The main module is realized using a MicroZed system-on-module (SoM) board. The MicroZed is a low-cost development board based on the Xilinx Zynq-7000 All Programmable SoC that integrates an ARM-based processor with a programmable FPGA. The SoM has two I/O headers that provide connection to two I/O banks on the programmable logic (PL). When the main module is plugged into a carrier card, the I/O headers are accessible in a way that is dependent on the carrier design.
@@ IV-A2 Converter Module
The data converter section of the converter module is based on LTC1740 6 Msps 14-bit ADC and LTC1668 50 Msps 16-bit DAC. A temperature-compensated crystal oscillator is present on the module to be used as the clock source for the ADC, DAC, and DSP blocks in the PL design. The main module power supply is implemented with high-efficiency switching regulators based on LTC3624. The design of the power supply for this module is challenging as most of the voltage regulators have switching frequencies that overlap with the spectrum commonly used for underwater acoustic communications. The resulting harmonic content of the switching noise causes poor performance, especially in multi-carrier communications. For these reasons, this board is implemented with a four-layer printed circuit board, analog and digital sections with isolated ground planes to keep high-speed digital traces away from analog signals, and to have separate power rails sourced from the power module for ADC and DAC.
@@ IV-A4 Power Module
The power module generates power rails from the battery supply. Supplies for data conversion modules and the preamplifier, are produced with low-noise linear voltage regulators. The system is protected against overvoltage, undervoltage, and short-circuit faults to maintain safe operation with high-capacity lithium-ion batteries. An external power switch can turn off or reboot the system without opening the enclosure. Both the power module and the signal module utilize card edge connectors used for the PCI Express bus standard which are ubiquitous, low-cost, and robust. The modules however have custom pinouts and form factors.
@@ IV-A5 Signal Module
The signal module has 98-pin board edge connectors similar to the power module. For digital I/O, 20 high-speed differential pin pairs are routed to the main module and separated from each other with ground connections. The analog I/O signals have four types of connections: preamplifier to ADC, DAC to power amplifier, transducer to preamplifier, and transducer to power amplifier. Each type of connection is replicated four times, forming four individual channels, and that at modules through jumpers.
@@
The assembled modules form a cylindrical structure that is long and wide, and it is shown in Fig. 5.
@@ IV-C Mechanical Implementation
The mechanical structure of the SEANet networking platform is shown in Fig. 7.
@@ V Experimental Results
In this section, we demonstrate our platform’s capabilities via various experiments. Initially, we assess power consumption under different operational settings, followed by data streaming and recording tests to validate both capabilities for generating datasets as well as validating hardware and software operations. We then establish an OFDM-based communication link, analyzing its performance across diverse modulations and bandwidths. Significantly, our platform’s OFDM physical layer and software-defined design achieve a data transfer rate of kbit/s, outperforming commercial modems. This demonstrates the platform’s adaptability in configuring data rates to suit application needs and reliability constraints. We also test a stop-and-wait ARQ protocol to showcase our data link layer capabilities. Lastly, the platform’s compatibility with the JANUS standard is demonstrated, indicating potential integration with existing JANUS-based systems. All tests are conducted in a marina setting (with a total depth of approximately 12 meters) using two prototype platforms deployed 5 meters apart at a depth of 3 meters (Fig. 8).
V-A Power Consumption
We evaluate the platform’s power consumption in different operational modes, directly by measuring from the battery terminals. The platform consume 5.97 W when idle and 6.11 W in receive mode. When transmitting, the platform consume from a minimum of 6.4 W to a maximum of 9.2 W. Considering the power consumed by the operations and the Ah capacity 4S lithium-ion battery mounted on each prototype, the expected life range goes from 44 hours when idle to 29 hours when in continuous transmission.
V-B Data Collection and Channel Characterization
@@ V-C Physical Layer
High Data Rate Communication Link. As previously mentioned, the software-defined design of the SEANet networking platform can be leveraged to optimize the link to achieve high data rates. For applications like video streaming, a high data rate device is paramount, however, none of the commercially available platforms is able to achieve data rates over kbit/s. For the following experiment, we used broadband transducers Neptune D/140, and Teledyne Reson TC4013 on the transmitter and receiver side respectively. The physical layer utilizes a kHz bandwidth on a kHz center frequency. For detection and synchronization, a maximum length sequence (m-sequence) preamble of 511 symbols is used with cross-correlation method. To reduce the packet duration and effective overhead, ms guard intervals are used along with OFDM blocks per frame. The OFDM blocks are loaded with differential QPSK modulated symbols over subcarriers, with a subcarrier bandwidth of Hz. The payload per packet is bytes and the raw data rate of this configuration is kbit/s. For forward error correction (FEC), we use a rate convolutional code with constraint length 12, and generator , along with termination. This channel coding, in conjunction with a random interleaver, yields a coded data rate of kbit/s, nearly triple the data rate achievable by commercial modems.
@@
We then solve the problem for , , , ms and target rates kbit/s and obtain for each rate. The solutions of the optimization problem for kbit/s and kbit/s are shown in the lower part of Fig. 12. The subcarriers in groups with are treated as null subcarriers, which also increases the power loaded to the rest of the subcarriers, potentially lowering the error rate for them. We repeat the experiment by using the calculated subcarrier assignments and measure the Packet Error Rate (PER) which is shown in Fig. 13. We showcase a large range of PER since for loss-tolerant applications we can use the higher data rates, for loss-intolerant applications we can choose lower data rates and lower corresponding PERs. For example, high data rates could support real-time video streaming, whereas control and telemetry streams would require a high degree of reliability.
V-D Link Layer
In this section, we demonstrate the potential of our proposed platform to establish reliable bidirectional links using link layer protocols. Specifically, we implement a stop-and-wait Automatic Repeat reQuest (ARQ) protocol [31]. Briefly, the protocol involves the sender transmitting variable-sized data packets and the receiver responding with ACKs or NACKs based on packet integrity. On receipt of a NACK or if neither ACK nor NACK is received within a pre-defined time (causing a timeout), the sender retransmits the packet. The physical layer, based on chirp spread spectrum, is identical for both forward and feedback links, with bandwidths of kHz and kHz respectively. To simulate increased packet loss, we apply dB attenuation to the transmitter.
@@ V-D Link Layer
V-E JANUS Integration
@@
In our demonstration, the toolkit processes I/Q samples in baseband mode with raw file format, and frequency conversion is performed on the modem. For practical purposes, we selected kHz center frequency to match our hardware specifications evaluated in Section V-B. The resulting configuration occupies larger frequencies than typical JANUS systems operating around kHz, which can be utilized to increase the data rate with as discussed in [33]. The data stream transmission is achieved by interfacing file input/output of the toolkit to the named pipes connected to DMA interfaces. In addition, significant portion of the signal up/down-conversion if offloaded to the PL by reducing the receiver’s sampling rate to ksps to decrease processing overhead and latency. Fig. 16 displays a received packet’s spectrogram with this setup, showing a discernible frequency hopping pattern Fig. 17 presents the JANUS toolkit output when a packet is received.

### v2

# Design and Performance Evaluation of SEANet, a Software-defined Networking Platform for the Internet of Underwater Things

@@ Abstract
This paper presents the design and performance evaluation of the SEANet platform, a software-defined acoustic modem designed for enhancing underwater networking and Internet of Underwater Things (IoUT) applications. Addressing the limitations of traditional acoustic modems, which suffer from low data rates and rigid architectures, SEANet introduces a versatile, adaptive framework capable of reconfiguring all layers of the protocol stack in real-time to accommodate diverse marine applications. The platform integrates high-performance, wideband data converters with modular hardware and software components, enabling real-time adaptation to changing environmental and operational conditions. Experimental evaluations conducted in oceanic settings demonstrate the SEANet capability to significantly exceed the performance of existing commercial underwater modems, supporting data rates up to 150 kbit/s and effectively doubling the performance metrics of conventional systems. Our robust testing also highlights the SEANet proficiency in channel estimation, Orthogonal Frequency-Division Multiplexing (OFDM) link establishment, and interoperability through the JANUS communication standard. Our results underscore the SEANet potential to transform underwater communication technologies, providing a scalable and efficient solution that supports high data rate applications and fosters the expansion of IoUT deployments.
@@
Acoustic modems have been shown to be viable solutions for long-range communications, as radio and optical technologies cannot be used by devices more than a few meters apart underwater [2]. Even if acoustic modems have been available for decades, current devices are beset by a host of limitations, including low data rates, cumbersome deployment procedures, and inflexible architectures. These constraints hinder the advancement of emerging underwater technologies that need real-time high-resolution data transmission [3, 4]. Common applications include transmitting imaging data through video streaming or side-scan sonar from autonomous underwater vehicles for monitoring and control purposes. These applications necessitate communication links with substantially higher data rates than those required for simpler sensing tasks. Currently, the data throughput capabilities of underwater networks are inadequate for high-quality data streaming, prompting investigations into potential solutions [5, 6]. Consequently, numerous IoUT applications are currently unfeasible with existing technology, motivating research and development of new devices designed to surmount these limitations.
@@ III-A2 Converter Module
The design of the converter module is motivated by the requirement to support diverse tasks in the underwater communications platform and to seamlessly integrate with various frontends. This necessitates the use of dedicated high-performance wideband data converters, which were previously unavailable as a main module option and are now incorporated within a specialized converter module. The converter module interfaces the main module with the rest of the modules. It has analog-to-digital (ADC) and digital-to-analog (DAC) data converters that provide analog interfaces to and from communication modules. In addition, it has voltage regulators for the main module, clock sources, and digital I/O ports. Since the platform is based on a direct conversion architecture and relies on digital mixers, to support wide range of frequencies used for underwater communications, fast data converters are chosen.
@@ III-A3 Communication Modules
The analog frontends of the platform must be tailored to suit specific applications, similar to commercial SDR platforms used in terrestrial networks. They should be engineered to meet performance criteria such as output power, bandwidth, and receiver sensitivity for the particular transducers they will utilize. This adaptability allows them to be configured for various tasks, including localization and acoustic sensing, and to function unidirectionally as either a beacon or a sensor. In this design, our focus is on wideband, high-frequency underwater acoustic communications applications.
@@ III-A4 Power Module
The platform design features a centralized power module, which supplies ready-to-use power rails to other modules, thereby streamlining the overall design process. This strategy not only optimizes the use of available board space for additional modules but also effectively separates noise-emitting components from critical signal pathways, enhancing the system’s overall performance. The power module is an interface to the battery unit and powers each module with different voltages using onboard voltage regulators. The module includes filters to reduce power supply noise and protection circuitry that continuously monitors the voltage and current levels delivered to the platform.
@@ III-A5 Signal Module
A separate signal module is included to provide expandability and support multi-channel operation. To provide this level of flexibility, a large number of signals need to be routed across different modules to create one-to-many connections. Using connectors on modules with cable assemblies is difficult to design and manage at this scale, so we use the following approach instead. The signal module routes digital and analog signals between the converter and communication modules and the transducer. As in the power module, each signal is accessible from any board in the stack. The signal module mounts reconfigurable board edge connectors to provide flexibility, e.g., it is possible to have more than one converter module attached to it. The signal module also has a connector for the transducer interface. This connection is driven by a duplexer circuit whose inputs can be switched between the power amplifier and preamplifier channels to allow a single transducer to be used for bidirectional communication. Even though transmit and receive chains can operate simultaneously, a half-duplex operation is implemented mainly due to the high cost of high-bandwidth acoustic transducers.
@@ IV Prototyping
This section outlines the implementation of the design introduced in Section III. We begin with the hardware implementation, then move on to the software components. Lastly, we address the development of the mechanical structure.
@@ IV-A1 Main Module
The main module executes the software architecture design while maintaining energy efficiency to operate on battery power. It should include all necessary interfaces for connectivity and sensors, although ports for user peripherals like displays or inputs are optional. Also, the module must be compact to fit within the size limitations of our underwater enclosure. We favor commercial computation modules due to their cost-effectiveness and the extensive community support available for both software and hardware development. The main module is realized using a MicroZed system-on-module (SoM) board. The MicroZed is a low-cost development board based on the Xilinx Zynq-7000 All Programmable SoC that integrates an ARM-based processor with a programmable FPGA. The SoM has two I/O headers that provide connection to two I/O banks on the programmable logic (PL). When the main module is plugged into a carrier card, the I/O headers are accessible in a way that is dependent on the carrier design.
@@ IV-A2 Converter Module
The core component of this module is the high-speed data converters that implement the direct conversion architecture. To reduce distortion from aliasing, these converters require high sampling rates exceeding the Nyquist rate of the ultrasonic frequencies utilized in acoustic communications. In addition to the large bandwidth requirement, they need to be high-precision to operate with low noise levels and have a good dynamic range to be able to receive transmissions from both nearby and far away transmitters. Most common data converter products for audio (high-precision, low-speed) or communication (high-speed, low-precision) applications do not meet these specific requirements. Therefore we have implemented this module as follows.
The data converter section of the converter module is based on LTC1740 6 Msps 14-bit ADC and LTC1668 50 Msps 16-bit DAC. A temperature-compensated crystal oscillator is present on the module to be used as the clock source for the ADC, DAC, and DSP blocks in the PL design. The power supply for the main module is implemented with high-efficiency switching regulators based on LTC3624. The design of the power supply for this module is challenging as most of the voltage regulators have switching frequencies that overlap with the spectrum commonly used for underwater acoustic communications. The resulting harmonic content of the switching noise causes poor performance, especially in multi-carrier communications. For these reasons, this board is implemented with a four-layer printed circuit board, analog and digital sections with isolated ground planes to keep high-speed digital traces away from analog signals, and to have separate power rails sourced from the power module for ADC and DAC.
@@ IV-A3 Communication Modules
The main difficulty in deploying the communication modules lies in interfacing with transducers. Standard amplifier configurations typically used in communications frequently face difficulties with these capacitive devices, leading to poor SNR performance. Although there are commercially available amplifiers designed for transducers operating in the ultrasonic frequency range, many are too large for our enclosure or require power inputs that do not match our platform specifications. Therefore, to achieve better integration, custom amplifier chains are utilized for these modules.
@@ IV-A4 Power Module
The modules described above need multiple voltage rails that differ in current requirements and noise specifications. Commercial power supply modules, while meeting the necessary voltage and current specifications, are unsuitable for two primary reasons. First, their noise characteristics at ultrasonic frequencies generate in-band noise, which impairs the SNR at the receiver. Second, they fail to conform to the mechanical specifications once assembled. Consequently, a custom power module has been developed that functions both as a structural element and a power source, deriving power rails directly from the battery supply.
The module utilizes low-noise linear voltage regulators to generate voltage rails for the data conversion modules and the preamplifier. The system is protected against overvoltage, undervoltage, and short-circuit faults to maintain safe operation with high-capacity lithium-ion batteries. An external power switch can turn off or reboot the system without opening the enclosure. Both the power module and the signal module utilize card edge connectors used for the PCI Express bus standard which are ubiquitous, low-cost, and robust. The modules however have custom pinouts and form factors.
@@ IV-A5 Signal Module
The signal module requires densely packed connectors for all analog and digital signal connections across the boards, while also accommodating future expansions with additional communication modules. The connectors must be durable and have a pitch that eases the hardware design of other modules. Although common coaxial connectors like SMA are sturdy, they necessitate substantial clearance for mating, which complicates the mechanical design. Smaller coaxial connectors, while less bulky, tend to detach easily and require labor-intensive inspection and maintenance. A more suitable approach involves adopting a unified single connector design that doubles as a mounting component for other modules. This design enables the signal module to also function as a structural element, securing the other modules in place. Our signal module has 98-pin board edge connectors similar to the power module. For digital I/O, 20 high-speed differential pin pairs are routed to the main module and separated from each other with ground connections. The analog I/O signals have four types of connections: preamplifier to ADC, DAC to power amplifier, transducer to preamplifier, and transducer to power amplifier. Each type of connection is replicated four times, forming four individual channels, and that at modules through jumpers.
@@
The assembled modules form a cylindrical structure that is long and wide (Fig. 5).
@@ IV-C Mechanical Implementation
For the mechanical implementation, our priority was to select an enclosure that met our design requirements. As enclosure size increases, its pressure rating typically decreases, which requires more robust construction to achieve the necessary pressure tolerance. While off-the-shelf enclosures made from materials like aluminum or titanium are available, they tend to be heavier, offer higher depth ratings, and are considerably more expensive. Although these enclosures might be suitable for specific applications, they add complexity, requiring specialized maintenance tools and additional buoyancy to offset their weight. Therefore, we decided on a lighter alternative for our enclosure to fulfill our design goals. The mechanical structure of the SEANet networking platform is shown in Fig. 7.
@@ V Experimental Results
In this section, we present an experimental evaluation of the proposed platform. Initially, we present results on characterization and verification of the platform’s functionality. Subsequently, we showcase potential applications across different domains, highlighting both its versatility and potential.
V-A Platform Verification
Initially, we assess power consumption under different operational settings, followed by data streaming and recording tests to validate both capabilities for generating datasets as well as validating hardware and software operations. We then establish an OFDM-based communication link, analyzing its performance across diverse modulations and bandwidths at different distances. All tests are conducted in a marina setting (with a total depth of approximately 12 meters) using two prototype platforms deployed 5 meters apart at a depth of 3 meters (Fig. 8).
V-A1 Power Consumption
We evaluate the platform’s power consumption in different operational modes, directly by measuring from the battery terminals. The platform consumes 5.97 W when idle and 6.11 W in receive mode. When transmitting, the platform consume from a minimum of 6.4 W to a maximum of 9.2 W. Considering the power consumed by the operations and the Ah capacity 4S lithium-ion battery mounted on each prototype, the expected life range goes from 44 hours when idle to 29 hours when in continuous transmission.
V-A2 Data Collection and Channel Characterization
@@ V-A3 Physical Layer
We conducted a longer distance test of our system, utilizing the same location but moving one of the platforms approximately m away. To improve the transmit response levels, we implemented a transformer-based matching circuit with the transducer, which, while extending the range, limits the available system bandwidth. At these distances, the frequency-dependent path loss, especially at the higher frequencies used by our platform, begins to present challenges.
As a result, we narrowed the bandwidth of our communication link to between kHz and kHz to manage these limitations. The results of BER vs. SNR from this test are depicted in Fig. 12. The data shows that system performance is affected by varying channel conditions and experiences notable fluctuations over time. In these tests, the peak SNR levels were lower than those in shorter-distance experiments, largely due to the limitations of the hardware, including communication modules and transducers. Additionally, the average results across different modulation and bandwidth settings varied, reflecting the changing instantaneous channel conditions. Nevertheless, the experiments demonstrate that by adjusting the physical layer parameters, satisfactory error performance can still be achieved at this range.
V-B Applications
In this section we showcase the variety of use cases and applications supported by our networking platform. First we establish a high data rate link with capable of supporting applications like underwater video streaming. Significantly, our platform’s OFDM physical layer and software-defined design achieve a data transfer rate of kbit/s, outperforming commercial modems. This demonstrates the platform’s adaptability in configuring data rates to suit application needs and reliability constraints. Then, we demonstrate end-to-end reliable transmission using our platform, emphasizing its utility as a network node that supports applications like file transfer. Lastly, we demonstrate the platform’s compatibility with heterogeneous networks by implementing the JANUS standard.
V-B1 High Data Rate Communication Link
For applications like video streaming, a high data rate device is paramount, however, none of the commercially available platforms is able to achieve data rates over kbit/s. As previously mentioned, the software-defined design of the SEANet networking platform can be leveraged to optimize the link to achieve high data rates. For the following experiment, we used broadband transducers Neptune D/140, and Teledyne Reson TC4013 on the transmitter and receiver side respectively. The physical layer utilizes a kHz bandwidth on a kHz center frequency. For detection and synchronization, a maximum length sequence (m-sequence) preamble of 511 symbols is used with cross-correlation method. To reduce the packet duration and effective overhead, ms guard intervals are used along with OFDM blocks per frame. The OFDM blocks are loaded with differential QPSK modulated symbols over subcarriers, with a subcarrier bandwidth of Hz. The payload per packet is bytes and the raw data rate of this configuration is kbit/s. For forward error correction (FEC), we use a rate convolutional code with constraint length 12, and generator , along with termination. This channel coding, in conjunction with a random interleaver, yields a coded data rate of kbit/s, nearly triple the data rate achievable by commercial modems.
@@ V-B1 High Data Rate Communication Link
We then solve the problem for , , , ms and target rates kbit/s and obtain for each rate. The solutions of the optimization problem for kbit/s and kbit/s are shown in the lower part of Fig. 13. The subcarriers in groups with are treated as null subcarriers, which also increases the power loaded to the rest of the subcarriers, potentially lowering the error rate for them. We repeat the experiment by using the calculated subcarrier assignments and measure the Packet Error Rate (PER) which is shown in Fig. 14. We showcase a large range of PER since for loss-tolerant applications we can use the higher data rates, for loss-intolerant applications we can choose lower data rates and lower corresponding PERs. For example, high data rates could support real-time video streaming, whereas control and telemetry streams would require a high degree of reliability as described in [6].
V-B2 End-to-end Reliable Transmission
Ensuring timely and reliable message delivery is crucial, particularly in scenarios where guaranteed delivery is essential to mitigate data loss due to channel conditions and interference [35]. For instance, offshore monitoring applications generate critical sensor data around infrastructures such as underwater pipelines or offshore oil platforms [36]. Although these systems often use wired connections for data transmission, exploiting the advantages of underwater acoustic networks necessitates the assurance of data reliability [37]. Moreover, this reliability is a fundamental requirement for enabling substantial data transfers, like file transfers, from underwater nodes.
In this section, we demonstrate the potential of our proposed platform to establish reliable bidirectional links using link layer protocols. Specifically, we implement a stop-and-wait Automatic Repeat reQuest (ARQ) protocol [38] to establish an end-to-end reliable communication link. Briefly, the protocol involves the sender transmitting variable-sized data packets and the receiver responding with ACKs or NACKs based on packet integrity. On receipt of a NACK or if neither ACK nor NACK is received within a pre-defined time (causing a timeout), the sender retransmits the packet. The physical layer, based on chirp spread spectrum, is identical for both forward and feedback links, with bandwidths of kHz and kHz respectively. To simulate increased packet loss, we apply dB attenuation to the transmitter.
@@ V-B2 End-to-end Reliable Transmission
V-B3 Heterogeneous Networks
Another application for our platform is its deployment in heterogeneous networks, which comprise a variety of vehicles, sensors, and buoys. These network nodes present diverse communication needs and deployment constraints, which restrict compatibility across different vendors platforms. Our platform acts as a conduit among these systems, enhancing the network by processing and forwarding packets among other underwater assets. To perform these functions, our platform’s software architecture is designed to emulate different waveforms and protocols. Furthermore, as highlighted in Section III-D, our platform supports a broad spectrum of frequencies. This feature allows it to communicate with various modems, thereby supporting interoperability between systems.
@@ V-B3 Heterogeneous Networks
In our demonstration, the toolkit processes I/Q samples in baseband mode with raw file format, and frequency conversion is performed on the modem. For practical purposes, we selected kHz center frequency to match our hardware specifications evaluated in Section V-A2. The resulting configuration occupies larger frequencies than typical JANUS systems operating around kHz, which can be utilized to increase the data rate with as discussed in [40].
The data stream transmission is achieved by interfacing file input/output of the toolkit to the named pipes connected to DMA interfaces. In addition, significant portion of the signal up/down-conversion if offloaded to the PL by reducing the receiver’s sampling rate to ksps to decrease processing overhead and latency. Fig. 17 displays a received packet’s spectrogram with this setup, showing a discernible frequency hopping pattern Fig. 18 presents the JANUS toolkit output when a packet is received.